OpenStudy (loser66):

which decimal digits occur as the final digit of a fourth power of an integer? Please, help

OpenStudy (fibonaccichick666):

well, the fifth happens to be the same as the original number's

OpenStudy (fibonaccichick666):

as for this, I guess we can find any fourth power last decimals by simply doing those on 1-9

OpenStudy (fibonaccichick666):

well, 0-9 but that's trivial

OpenStudy (loser66):

Hey, is there any other method than just list the digit out?

OpenStudy (fibonaccichick666):

uhm, there is, but uh, number theory, uh I got around using the formulas by doing this last time, so uh, I don't remember :/

OpenStudy (loser66):

What does it mean by saying: the last digit of a^2 is a^2mod10?

ganeshie8 (ganeshie8):

whats the remainder when you divide 2315 by 10 ?

OpenStudy (loser66):

5

OpenStudy (loser66):

I think I got it.

ganeshie8 (ganeshie8):

Hey sorry there was some issue with electricity last night

OpenStudy (loser66):

I know and after sleeping, I figured it out :)

ganeshie8 (ganeshie8):

nice :) so what have you figrued out ``` What does it mean by saying: the last digit of a^2 is a^2mod10? ```

OpenStudy (loser66):

all of them :) However, I have another question on this topic. Let me post it here.

OpenStudy (loser66):

If \(a^k \equiv b^k (mod m)\\a^{k+1}\equiv b^{k+1}(modm)\) \(a, b, m, k \in \mathbb Z; m>0, k>0, (a,m)=1\) Prove \(a\equiv b(modm)\)

OpenStudy (loser66):

My question: Why can't we go directly from the problem to get the answer? Since we have \(a^k \equiv b^k (mod m)\) and \(a^{k+1}= a^k *a\) Hence \(a*a^k \equiv b*b^k \implies a\equiv b\)

ganeshie8 (ganeshie8):

are we given that \(a^k \equiv b^k \pmod{}\) is valid for all \(k\gt 0\) ?

OpenStudy (loser66):

Why do we have to expand \(a^{k+1}-b^{k+1}= (a-b)(a^{k} +a^{k-1}b +\cdots +b^{k})\) and then use the fact that (a, m) =1 to show that the LHS divided by m iff m | (a-b) to get \(a\equiv b (mod m)\)

OpenStudy (loser66):

Yes

ganeshie8 (ganeshie8):

then the question makes no sense, why can't we directly plugin \(k=1\) ?

ganeshie8 (ganeshie8):

since we are given that \(a^k \equiv b^k \pmod{m}\) is valid for all \(k\gt 0\) plugging in \(k=1\) gives \(a^1 \equiv b^1 \pmod{m}\)

OpenStudy (loser66):

We cannot, since k is arbitrary,

ganeshie8 (ganeshie8):

you have said the given statemetn is true for all \(k\gt 0\) ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

or do you mean \(k\) is some "fixed" positive integer ?

ganeshie8 (ganeshie8):

could you take a screenshot of the problem and post ?

OpenStudy (loser66):

For all k > 0

OpenStudy (loser66):

sure

OpenStudy (loser66):

Problem 26

ganeshie8 (ganeshie8):

Okay, we should interpret \(k\) as some specific "fixed" positive integer

OpenStudy (loser66):

ok, next?

ganeshie8 (ganeshie8):

My question: Why can't we go directly from the problem to get the answer? Since we have \(a^k \equiv b^k (mod m)\) and \(a^{k+1}= a^k *a\) Hence \(a*a^k \equiv b*b^k \implies a\equiv b\) How did you get that ?

OpenStudy (loser66):

oh, yeah, it is invalid!! we have \(a\equiv b \\c\equiv d \\ac \equiv bd\) not backward.

ganeshie8 (ganeshie8):

it is perfectly fine, you need to add proper justifications, thats all

ganeshie8 (ganeshie8):

\(a^k\equiv b^k \pmod{m}\) and \(a^{k+1}\equiv b^{k+1}\pmod{m}\) together imlies \(a*a^k \equiv b*b^k \equiv b*a^k \pmod{m}\)

ganeshie8 (ganeshie8):

so we have \(a*a^k \equiv b*a^k \pmod{m}\) nothing fancy, fine so far, right ?

OpenStudy (loser66):

Yes

ganeshie8 (ganeshie8):

but what exactly does that mean

OpenStudy (loser66):

\(a\equiv b\)

ganeshie8 (ganeshie8):

\(a*a^k \equiv b*a^k \pmod{m} \iff m \mid (a*a^k - b*a^k)\)

ganeshie8 (ganeshie8):

factor out \(a^k\) and apply Euclid lemma

OpenStudy (loser66):

and factor a^k out and use the fact that m not | a to conclude that m | (a-b)

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (loser66):

Thank you so much.

ganeshie8 (ganeshie8):

np